OpenStudy (anonymous):

Let T : C∞(R) −→ C∞(R) be the map given by T(f) = f′′− 2f′− 3f . Find a linear combination a sin(x) + b cos(x) which solves the equation f′′− 2f′− 3f = 2 cos(x).

OpenStudy (anonymous):

\[f(x)=c_1e^{2x}+c_2+\frac{1}{10}(-15x-8\sin(x)-4\cos(x))\]

OpenStudy (anonymous):

how did you get that!!!!!

OpenStudy (anonymous):

I solved the ODE you posted.

OpenStudy (anonymous):

and how is that equal to 2cosx?

OpenStudy (anonymous):

If you take that function and evaluate\[f''-2f'-3f,\]you will get\[2\cos(x).\]

OpenStudy (anonymous):

hmm ok